# Principal component analysis balancing prediction and approximation accuracy for spatial data

### v1

@@ 5.2 Analysis of HER2-Positive Breast Tumor Spatial Transcriptomics
We analyze the human epidermal growth factor receptor 2 positive (HER2+) breast tumor data, which includes expression measures of genes in HER2+ tumors from eight individuals (Andersson et al., 2021). We focus on the first tumor section of the last individual, coded as sample H1 in Andersson et al. (2021), where the data consists of expression counts of $15,030$ genes on $613$ tissue locations. Following the same filtering steps as in Shang and Zhou (2022), we removed genes with non-zero expression at less than 20 locations and those confounded by technical artifacts (Andersson et al., 2021), as well as locations with non-zero expressions for less than 20 genes. The filtered set of data contains $10,053$ genes measured on $607$ spots, which is then normalized via regularized negative binomial regression as implemented by the Seurat R package (Hafemeister and Satija, 2019).
Due to the large number of genes along with the noise present in the expression measurements (Rao et al., 2021), dimension reduction is commonly done before the main analytic tasks on genomic data to extract biologically meaningful signals as well as to facilitate computation; see Sun et al. (2019); Zhao et al. (2021); Shang and Zhou (2022) for related examples. We focus on two modeling tasks: spatial extrapolation and domain detection. In addition to the three PCA methods (PCA, predictive PCA, RapPCA) that were investigated in previous sections, we also include spatial PCA for comparison, which is a probabilistic PCA algorithm for spatial transcriptomics data incorporating spatial proximity information in the modeling of PC scores (Shang and Zhou, 2022).
The first application is more similar to the analysis in Section 5.1, where we are interested in predicting the gene expression PC scores at tissue locations with no measurements. This prediction problem is involved, for example, in the reconstruction of high-resolution spatial maps of gene expression (Gryglewski et al., 2018; Shang and Zhou, 2022). For each dimension reduction method, we extract the top 20 PCs from the expression of all genes, and conduct 10-fold cross-validation to assess TMSE, MSPE and MSRE-trn, where spatial prediction is done by random forest followed by TPRS smoothing as in previous sections. For spatial PCA, we also include results from its built-in spatial prediction function (from the SpatialPCA R package) for completeness. Table 2 summarizes the overall metrics along with individual prediction MSEs for the top 3 PCs. RapPCA demonstrates significant advantage in overall prediction accuracy (MSPE) compared to all other methods, followed by predictive PCA which also has the lowest per-PC MSEs for the top 3 PCs. Also, consistent with findings from previous sections, RapPCA achieves the optimal trade-off between prediction and approximation gaps, as reflected by TMSE.
@@
Figure 5 shows the top 3 PC scores from the expressions of all genes. We observe that spatial PCA and predictive PCA produce spatially smoother surfaces of PC scores compared to the other two methods. This is natural for predictive PCA since it prioritizes spatial predictability of the PC scores which results in stronger spatial smoothing; while for spatial PCA, the prediction-representation balance is implicit and less straightforward to interpret. The gap in prediction performance comparing spatial or predictive PCA with RapPCA may be the consequence of over-smoothing, which also indicates the effectiveness of explicit optimization for the prediction-representation balance achieved by (3). Figure 11 in Appendix B.2 visualizes the smoothed high-resolution maps of PC scores based on predictions at unmeasured locations.
Figure 5: Top 3 PC scores based on gene expression in the HER2-positive breast tumor data
As a second example, we investigate the problem of domain detection following the dimension reduction step. Domain detection is the task where different sections of tissues are identified as various spatially coherent and functionally distinct regions (Dong and Zhang, 2022; Shang and Zhou, 2022), providing helpful insights into the biological function of tissues. To this end, we conduct domain detection via walktrap clustering algorithm (Pons and Latapy, 2005) using the top 20 PCs extracted by each variant of PCA. The number of clusters is set to align with the “ground truth” labels based on pathologist annotations of tissue regions in the original study (Andersson et al., 2021). The hyperparameters $\gamma,\lambda_{1},\lambda_{2}$ are chosen based on silhouette value measuring the difference between nearest-cluster versus intra-cluster distances.
Figure 6 visualizes the detected spatial domains based on PCs extracted from different dimension reduction procedures, compared to the ground truth labels. In general, the relative performance of each approach differs across spatial domains. For example, PCA fails to detect the adipose tissue region and produces noisier results for regions surrounding invasive cancer cells. Predictive PCA and RapPCA both mis-classify part of the cancer in situ region as invasive cancer, where predictive PCA results are noisier for the top left region, and RapPCA shows larger uncertainty regarding adipose tissue. Spatial PCA, on the other hand, infers part of the connective tissue to be cancerous.
@@
To make a more in-depth comparison, we examine the breakdown of clustering accuracy by ground truth labels, in other words, whether or not each approach correctly classifies the spots belonging to each spatial domain, with the undetermined cluster removed from comparison. Figure 7 presents this breakdown for each method in terms of precision, recall and F1 score as a composite measure. As observed from the spatial domain plots, the relative accuracy for each method differs by region, and in particular, RapPCA has the best performance in both precision and recall on the identification of invasive cancer regions, whereas spatial PCA shows an advantage in recall for cancer in situ.
Figure 7: Breakdown of domain detection accuracy by true label. Note that the metrics for the adipose tissue region are not well-defined for PCA because it fails to detect any spot in this region.
@@ 6 Discussion
We discussed how existing dimension reduction algorithms fit into this general framework of optimizing either criterion, and proposed a flexible interpolation between them that achieves the optimal representability-predictability trade-off. Our proposal, called RapPCA, can also handle high-dimensional covariates and non-linear relationships. Simulation studies under different scenarios illustrate the gain in downstream prediction accuracy by our method, which also achieves smaller overall errors when both predictability and representability are taken into account. Applications to real datasets in different scientific areas further demonstrate the utility of our proposed method, even in analytic tasks that do not explicitly involve prediction.
@@
While we used TMSE to guide the choice of tuning parameters in our examples, in practice tuning parameter selection could be driven by the specific analytic goal in practice. Researchers could examine the trend of prediction, representation and total errors by each tuning parameter, and in particular $\gamma$ in (3), on a set of test data, and choose the combination leading to a desired trade-off. Our empirical evaluations suggest that striking such balance with our proposed method would most often improve, or at least maintain similar, overall errors compared to existing alternatives.
The procedure described in Algorithm 1 depends on a specified number $r$ of PCs. Since increasing $r$ will reduce the total MSRE and increase the total MSPE, one could examine the trajectory of MSPE and MSRE with respect to $r$, and identify the elbow of these curves as a practical choice of $r$; this reflects the threshold beyond which adding more PCs would not alter either the prediction or representation meaningfully. Practitioners may also be interested in estimating the variability of evaluation metrics such as TMSE, MSPE and MSRE to better understand the trade-off between them. While our method does not directly provide uncertainty estimates, one could conduct bootstrap to assess this, with the potential caveat of sample splitting on correlated samples. Though Theorem 1 establishes an exact solution to the optimization problem (3) that, in theory, does not contain numerical errors, computational inaccuracy may occur in scenarios with large sample size or high dimensionality, since the quality of eigen-decomposition could suffer when the condition number of the matrix involved is large. Alternative numerical techniques could improve the accuracy and stability of the proposed dimension reduction procedure.
@@ Appendix A Proof
The superscript $(l)$ will be suppressed for simplicity. We denote $y=Y^{(l)}$ and start by examining the first term in the objective function, which, under the constraint, can be expressed as
$\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}$ $=\text{tr}\left((y-uv^{\top})^{\top}(y-uv^{\top})\right)$
@@
where $\text{tr}(\cdot)$ denotes the trace of a matrix, $\propto$ means equal up to a constant not depending on $u,v,\alpha$ or $\beta$, and we have made use of the cyclic property of trace. Since $v^{\top}v=1$ and ${\mathcal{S}}^{\top}{\mathcal{S}}=I$, denoting $q:={\mathcal{T}}^{\top}v$, (4) continues as
$\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}$ $=-2\text{tr}\left(q^{\top}\mathcal{D}^{2}q\right)+\text{tr}\left(q^{\top}\mathcal{D}^{2}q\right)=-{q^{\top}\mathcal{D}^{2}q}.$
@@ B.1 Optimality of Proposed Solution
We explore different values of tuning parameters $\gamma,\lambda_{1},\lambda_{2}$ and compare the objective function $f_{\gamma,\lambda_{1},\lambda_{2}}(u,v,\eta)$ evaluated at the optimal solution $(\tilde{u},\tilde{v},\tilde{\eta})$ versus the modified solution $(\tilde{u},v^{*}(\theta),\tilde{\eta})$ for $\theta\in[0,2\pi)$. We plot their differences against $\theta$ for various values of tuning parameters in Figures 8 and 9. Note that the optimal $\theta$, i.e., the value leading to the optimal loading $v^{*}(\theta)$, may differ with different sets of tuning parameters $\gamma,\lambda_{1}$ and $\lambda_{2}$. The fact that each curve is always above or equal to zero, and reaches zero exactly once, verifies the uniqueness of the optimal solution, as well as the optimality of the proposed form of $\tilde{v}$, at least for fixed $\tilde{u}$ and $\tilde{\eta}$.
@@
Figure 8: Difference between the objective function evaluated over a range of values for the first two parameters in the loading vector and the optimum achieved by our algorithm. The $x$-axis is parameterized by $\theta$, the polar coordinate angular representation of the candidate parameters, and we fix $\lambda_{1}=\lambda_{2}$. Each panel corresponds to a value of $\lambda_{1}$ and $\lambda_{2}$.
@@
Figure 9: Difference between the objective function evaluated over a range of values for the first two parameters in the loading vector and the optimum achieved by our algorithm. The $x$-axis is parameterized by $\theta$, the polar coordinate angular representation of the candidate parameters, and we fix $\lambda_{1}=1$. Each panel corresponds to a value of the ratio $\lambda_{2}/\lambda_{1}$.
B.2 Additional Simulation and Data Analysis Results
@@
Figure 10 is an analog to Figure 1 presented in our simulation studies, which visualizes the variation of TMSE, MSPE and MSRE-trn for the first PC with respect to the combination of $\lambda_{1}$ and $\lambda_{2}$. For each combination, the additional tuning parameter $\gamma$ is chosen such that it leads to the lowest TMSE. In Scenarios 1 through 3, the median values of the selected parameters $(\lambda_{1},\lambda_{2}/\lambda_{1})$ via cross-validation are (0.5, 1), (1, 0.75) and (0.5, 0.25) respectively, all of which are reasonably close to near-optimal region(s) in Figure 10.
@@
Figure 10: Average TMSE, MSPE and MSRE-trn for the first PC arcoss 100 replicates of data, plotted against the combination of $\lambda_{1}$ and the ratio $\lambda_{2}/\lambda_{1}$.
@@
Figure 11 presents the smoothed high-resolution maps of predicted PC scores, obtained by each dimension reduction approach. We observe that predictive PCA produces the smoothest predicted map due to its emphasis on the predictability of PC scores; however, it could over-smooth and omit meaningful spatial variations in gene expression, as can be seen in the high-resolution map for the third PC. For spatial PCA, while the extracted low-resolution PCs also demonstrate smoothness as predictive PCA (see Figure 5), they are not guaranteed to be well predictable and the extrapolated high-resolution map appears noisier. RapPCA achieves a reasonable balance among all algorithms.
@@
Figure 11: Smoothed PC scores based on predicted high-resolution gene expression maps in the HER2-positive breast tumor data

### v2

@@ 5.2 Analysis of HER2-Positive Breast Tumor Spatial Transcriptomics
We analyze the human epidermal growth factor receptor 2 positive (HER2+) breast tumor data, which includes expression measures of genes in HER2+ tumors from eight individuals (Andersson et al., 2021). We focus on the first tumor section of the last individual, coded as sample H1 in Andersson et al. (2021), where the data consists of expression counts of $15,030$ genes on $613$ tissue locations. Following the same filtering steps as in Shang and Zhou (2022), we removed genes with non-zero expression at less than 20 locations and those confounded by technical artifacts (Andersson et al., 2021), as well as locations with non-zero expressions for less than 20 genes. The filtered set of data contains $p=10,053$ genes measured on $n=607$ spots, which is then normalized via regularized negative binomial regression as implemented by the Seurat R package (Hafemeister and Satija, 2019).
Due to the large number of genes along with the noise present in the expression measurements (Rao et al., 2021), dimension reduction is commonly performed before the main analytic tasks on genomic data to extract biologically meaningful signals as well as to facilitate computation; see Sun et al. (2019); Zhao et al. (2021); Shang and Zhou (2022) for related examples. We focus on two modeling tasks: spatial extrapolation and domain detection. In addition to the three PCA methods (PCA, predictive PCA, RapPCA) that were investigated in previous sections, we also include spatial PCA in our comparison. Spatial PCA is a probabilistic PCA algorithm for spatial transcriptomics data which incorporates spatial proximity information when modeling PC scores (Shang and Zhou, 2022).
The first application is more similar to the analysis in Section 5.1, where we are interested in predicting the gene expression PC scores at tissue locations with no measurements. This prediction problem is involved, for example, in the reconstruction of high-resolution spatial maps of gene expression (Gryglewski et al., 2018; Shang and Zhou, 2022). For each dimension reduction method, we extract the top 20 PCs from the expression of all genes, and conduct 10-fold cross-validation to assess TMSE, MSPE and MSRE-trn, where spatial prediction is done by random forest followed by TPRS smoothing as in previous sections. For spatial PCA, we also include results from its built-in spatial prediction function (using the SpatialPCA R package) for completeness. Table 2 summarizes the overall metrics along with individual prediction MSEs for the top 3 PCs. RapPCA demonstrates significant advantage in overall prediction accuracy (MSPE) compared to all other methods, followed by predictive PCA which also has the lowest per-PC MSEs for the top 3 PCs. Also, consistent with findings from previous sections, RapPCA achieves the optimal trade-off between prediction and approximation gaps, as reflected by TMSE.
@@
Figure 4 visualizes the top 3 PC scores from the expressions of all genes. We observe that spatial PCA and predictive PCA produce spatially smoother surfaces of PC scores compared to the other two methods. This is natural for predictive PCA since it prioritizes spatial predictability of the PC scores, which results in stronger spatial smoothing; for spatial PCA, the prediction-representation balance is implicit and less straightforward to interpret. The gap in the prediction performance of spatial or predictive PCA compared with RapPCA may be the consequence of over-smoothing, which also indicates the effectiveness of explicit optimization for the prediction-representation balance achieved by (4). Figure S5 in Appendix B.3 visualizes the smoothed high-resolution maps of PC scores based on predictions at unmeasured locations.
Figure 4: Spatial distribution of the top 3 PC score values extracted by PCA, predictive PCA, RapPCA and spatial PCA, based on measured gene expression in the HER2-positive breast tumor data.
In the second application, we investigate the problem of domain detection following the dimension reduction step. In domain detection, different sections of tissues are identified as various spatially coherent and functionally distinct regions (Dong and Zhang, 2022; Shang and Zhou, 2022), providing helpful insights into the biological function of tissues. To this end, we conduct domain detection via the walktrap clustering algorithm (Pons and Latapy, 2005) using the top 20 PCs extracted by each algorithm. The number of clusters is set to align with the “ground truth” labels based on pathologist annotations of tissue regions in the original study (Andersson et al., 2021). The hyperparameters $\gamma,\lambda_{1},\lambda_{2}$ are chosen based on silhouette value measuring the difference between nearest-cluster versus intra-cluster distances (Rousseeuw, 1987).
Figure 5 visualizes the detected spatial domains based on PCs extracted from different dimension reduction procedures, compared to the ground truth labels. In general, the relative performance of each algorithm differs across spatial domains. For example, PCA fails to detect the adipose tissue region and produces noisier results for regions surrounding invasive cancer cells. Predictive PCA and RapPCA both mis-classify part of the cancer in situ region as invasive cancer, where predictive PCA results are noisier for the top left region, and RapPCA shows larger uncertainty regarding adipose tissue. Spatial PCA, on the other hand, infers part of the connective tissue to be cancerous.
@@
To make a more in-depth comparison, we examine the breakdown of clustering accuracy by ground truth labels; that is, whether or not each approach correctly classifies the spots belonging to each spatial domain, with the undetermined cluster removed from comparison. Figure 6 presents this breakdown for each method in terms of precision, recall and F1 score. As observed from the spatial domain plots, the relative accuracy for each method differs by region, and in particular, RapPCA has the best performance in both precision and recall on the identification of invasive cancer regions, whereas spatial PCA shows an advantage in recall for cancer in situ.
Figure 6: Breakdown of domain detection accuracy by true label for each algorithm. Note that the metrics for the adipose tissue region (bottom row) are not well-defined for PCA because it fails to detect any spot in this region.
@@ 6 Discussion
We discussed how existing dimension reduction algorithms fit into this general framework of optimizing either criterion, and proposed a flexible interpolation between them that achieves the optimal representability-predictability trade-off. Our proposal, called RapPCA, can also handle high-dimensional predictors (including covariates and spatial basis) and non-linear relationships between covariates and PC scores. Simulation studies under different scenarios illustrate the gain in downstream prediction accuracy by our method, which also achieves smaller overall errors when both predictability and representability are taken into account. Applications to real datasets in different scientific domains further demonstrate the utility of our proposed method, even in analytic tasks that do not explicitly involve prediction.
@@
While we used TMSE to guide the choice of tuning parameters in most of our examples, in practice tuning parameter selection could be driven by the specific analytic goal. For example, we used silhouette value reflecting the goodness of clustering for the domain detection task in Section 5.2; while for the health effect analysis of air pollution in Jandarov et al. (2017), the main focus was spatial extrapolation accuracy and hence prediction error is of primary interest. Researchers could examine the trend of prediction, representation and total errors by each tuning parameter, and in particular $\gamma$ in (4), on a set of test data, and choose the combination leading to the desired trade-off. Our empirical evaluations suggest that striking such balance with our proposed method would most often improve, or at least maintain similar, overall errors compared to the existing alternatives.
The procedure described in Algorithm 1 depends on a specified number $r$ of PCs. Since increasing $r$ will reduce the total MSRE and increase the total MSPE, one could examine the trajectory of MSPE and MSRE with respect to $r$, and identify the ‘elbow’ of these curves as a practical choice of $r$; this reflects the threshold beyond which adding more PCs would not alter either the prediction or representation meaningfully. Practitioners may also be interested in estimating the variability of evaluation metrics such as TMSE, MSPE and MSRE to better understand the trade-off between them. While our method does not directly provide uncertainty estimates, one could conduct bootstrap to assess this, with the potential caveat of sample splitting on correlated data.
@@ Appendix A Proof
Suppressing the superscript $(l)$ for simplicity, we denote $y=Y^{(l)}$ and start by examining the first term in the objective function, which can be expressed as
$\left\lVert Y^{(l)}-Y^{(l)}vv^{\top}\right\rVert_{F}^{2}$ $=\text{tr}\left((y-yvv^{\top})^{\top}(y-yvv^{\top})\right)$
@@
where $\text{tr}(\cdot)$ denotes the trace of a matrix, $\propto$ means equal up to a constant not depending on $v,\alpha$ or $\beta$, and we have made use of the cyclic property of trace. Since $v^{\top}v=1$ and ${\mathcal{S}}^{\top}{\mathcal{S}}=I$, denoting $q:={\mathcal{T}}^{\top}v$, (6) can be written as
$\left\lVert Y^{(l)}-Y^{(l)}vv^{\top}\right\rVert_{F}^{2}$ $=-2\text{tr}\left(q^{\top}\mathcal{D}^{2}q\right)+\text{tr}\left(q^{\top}\mathcal{D}^{2}q\right)=-{q^{\top}\mathcal{D}^{2}q}.$
@@ B.1 Optimality of Proposed Solution
We explore different values of tuning parameters $\gamma,\lambda_{1},\lambda_{2}$ and compare the objective function $f_{\gamma,\lambda_{1},\lambda_{2}}(u,v,\eta)$ evaluated at the optimal solution $(\tilde{u},\tilde{v},\tilde{\eta})$ versus the modified solution $(\tilde{u},v^{*}(\theta),\tilde{\eta})$ for $\theta\in[0,2\pi)$. We plot their differences against $\theta$ for various values of tuning parameters in Figures S1 and S2. Note that the optimal $\theta$, i.e., the value leading to the optimal loading $v^{*}(\theta)$, may differ with different sets of tuning parameters $\gamma,\lambda_{1}$ and $\lambda_{2}$. The fact that each curve is always above or equal to zero, and reaches zero exactly once, verifies the uniqueness of the optimal solution, as well as the optimality of the proposed form of $\tilde{v}$, at least for fixed $\tilde{u}$ and $\tilde{\eta}$.
@@
Figure S1: Difference between the objective function evaluated over a range of values for the first two parameters in the loading vector and the optimum achieved by our algorithm. The $x$-axis is parameterized by $\theta$, the polar coordinate angular representation of the candidate parameters, and we fix $\lambda_{1}=\lambda_{2}$. Each panel corresponds to a value of $\lambda_{1}$ and $\lambda_{2}$.
@@
Figure S2: Difference between the objective function evaluated over a range of values for the first two parameters in the loading vector and the optimum achieved by our algorithm. The $x$-axis is parameterized by $\theta$, the polar coordinate angular representation of the candidate parameters, and we fix $\lambda_{1}=1$. Each panel corresponds to a value of the ratio $\lambda_{2}/\lambda_{1}$.
B.2 Additional Simulation Results
@@
Figure S3 is an analog of Figure 1 presented in our simulation studies, which visualizes the variation of TMSE, MSPE and MSRE-trn for the first PC with respect to the combination of $\lambda_{1}$ and $\lambda_{2}$, instead of $\gamma$. For each combination, the additional tuning parameter $\gamma$ is chosen such that it leads to the lowest TMSE. In Scenarios 1 through 3, the median values of the selected parameters $(\lambda_{1},\lambda_{2}/\lambda_{1})$ via cross-validation are (0.5, 1), (1, 0.75) and (0.5, 0.25) respectively, all of which are reasonably close to near-optimal region(s) in Figure S3. This reflects the effectiveness of the cross-validation procedure to control the degree of freedom of coefficients $\alpha$ and $\beta$ in (4).
B.3 Additional Data Analysis Results
@@
Figure S3: Average TMSE, MSPE and MSRE-trn for the first PC extracted by RapPCA arcoss 100 replicates of data, plotted against the combination of $\lambda_{1}$ and the ratio $\lambda_{2}/\lambda_{1}$.
For the TRAP data analysis presented in Section 5.1, Figure S4 presents the smoothed, finer-grain PC scores obtained by each method across the study region. We observe similar spatial patterns in the distribution of the PC scores across different methods, except for the south end of the study region for the third PC where PCA identifies a stronger signal than RapPCA and predictive PCA. In particular, all methods highlight regions near the airport and/or around major roads (the dark red area) for the first PC, indicating aircraft and road traffic emissions as a major source of overall pollution level. This is consistent with the large contributions of BC as well as UFP with small or moderate sizes (Zhang et al., 2019; Bendtsen et al., 2021), as reflected by the loadings of UFP (with relatively small particle sizes) and BC in Figure 3.
Figure S4: Smoothed PC scores of pollutant concentrations from the Seattle TRAP data, based on spatial extrapolation following each dimension reduction algorithm.
@@
For the spatial transcriptomics application in Section 5.2, Figure S5 presents the smoothed high-resolution maps of predicted PC scores, obtained by spatial prediction following each dimension reduction approach. We observe that predictive PCA produces the smoothest predicted map due to its emphasis on the predictability of PC scores; however, it could over-smooth and omit meaningful spatial variations in gene expression, as can be seen in the high-resolution map for the third PC. For spatial PCA, while the extracted low-resolution PCs also demonstrate smoothness as predictive PCA (see Figure 4), they are not guaranteed to be well predictable and the extrapolated high-resolution map appears noisier. RapPCA achieves a reasonable balance among all algorithms.
@@
Figure S5: Smoothed PC scores based on predicted high-resolution gene expression maps following each dimension reduction algorithm in the HER2-positive breast tumor data.